# Arens extensions of disjointness preserving multilinear operators on Riesz spaces

### v1

@@ 2 Disjointness preserving linear and multilinear operators
We conclude this preparatory section by proving that every disjointness preserving positive multilinear form can be written as the product of disjointness preserving positive linear functionals.
Theorem 2.6.
Let $E_{1},\ldots,E_{m}$ be Riesz spaces and let $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow\mathbb{R}$ be a positive $m$-linear form. If $A$ is disjointness preserving, then there exist disjointness preserving positive linear functionals $\varphi_{1}\in E_{1}^{\sim},\ldots,\varphi_{m}\in E_{m}^{\sim}$ such that $A(x_{1},\ldots,x_{m})=\varphi_{1}(x_{1})\cdots\varphi_{m}(x_{m})$ for all $x_{1}\in E_{1},\ldots,x_{m}\in E_{m}$.
Proof.
We proceed by induction on $m$. The result is immediate for $m=1$. Now we assume that the result holds for $m=k$ and we check its validity for $m=k+1$. To do so, let a positive disjointness preserving $(k+1)$-linear form $A\colon E_{1}\times\cdots\times E_{k+1}\longrightarrow\mathbb{R}$ be given. Assuming, without loss of generality, that $A\neq 0$, let $x_{0}^{1}\in E_{1}^{+},\ldots,x_{0}^{k}\in E_{k}^{+}$ be such that the map
$A_{x_{0}^{1},\ldots,x_{0}^{k}}\colon E_{k+1}\longrightarrow\mathbb{R}~{},~{}A_{x_{0}^{1},\ldots,x_{0}^{k}}(x_{k+1})=A(x_{0}^{1},\ldots,x_{0}^{k},x_{k+1}),$
is a non-zero disjointness preserving positive linear operator. Note that $A_{x_{0}^{1},\ldots,x_{0}^{k}}$ is a Riesz homomorphism by the proof of Theorem 2.1(1). And it is easy to check that
$A_{x_{1}+x_{0}^{1},\ldots,x_{k}+x_{0}^{k}}\geq A_{x_{0}^{1},\ldots,x_{0}^{k}}\geq 0\quad\text{and}\quad A_{x_{1}+x_{0}^{1},\ldots,x_{k}+x_{0}^{k}}\geq A_{x_{1},\ldots,x_{k}}\geq 0$
for all $x_{1}\in E_{1}^{+},\ldots,x_{k}\in E_{k}^{+}$. Applying Kutateladze’s Theorem [5, Theorem 2.50] for the inequalities above, there exist positive real numbers $t$ and $s$ such that
$A_{x_{0}^{1},\ldots,x_{0}^{k}}=tA_{x_{1}+x_{0}^{1},\ldots,x_{k}+x_{0}^{k}}\quad\text{and}\quad A_{x_{1},\ldots,x_{k}}=sA_{x_{1}+x_{0}^{1},\ldots,x_{k}+x_{0}^{k}}.$
As $A_{x_{0}^{1},\ldots,x_{0}^{k}}\neq 0$, $t\neq 0$, hence $A_{x_{1},\ldots,x_{k}}=\dfrac{s}{t}\,A_{x_{0}^{1},\ldots,x_{0}^{k}}$. This shows that, for all $x_{1}\in E_{1}^{+},\ldots,x_{k}\in E_{k}^{+}$, there exists $\gamma(x_{1},\ldots,x_{k})\in\mathbb{R}^{+}$ such that
$A_{x_{1},\ldots,x_{k}}=\gamma(x_{1},\ldots,x_{k})A_{x_{0}^{1},\ldots,x_{0}^{k}}.$
Taking $x_{0}^{k+1}\in E^{+}$ such that $A_{x_{0}^{1},\ldots,x_{0}^{k}}(x_{0}^{k+1})=1$, we get
$\gamma(x_{1},\dots,x_{k})=A_{x_{1},\dots,x_{k}}(x_{0}^{k+1})=A(x_{1},\dots,x_{k},x_{0}^{k+1})\text{~{}for all~{}}x_{1}\in E_{1}^{+},\dots,x_{k}\in E_{k}^{+}.$
Thus, the map $\gamma\colon E_{1}^{+}\times\cdots\times E_{k}^{+}\to\mathbb{R}^{+}$ is additive in each variable, and by [17, Theorem 2.3] it extends uniquely to a positive $k$-linear mapping $\widetilde{\gamma}\colon E_{1}\times\cdots\times E_{k}\to\mathbb{R}$ given by
$\widetilde{\gamma}(x_{1},\ldots,x_{k})=\sum_{{n_{i}=0,1}\atop{1\leq i\leq k}}(-1)^{\sum n_{i}}\gamma(x_{1n_{1}},\ldots,x_{kn_{k}}),$
where $x_{j}=x_{j0}-x_{j1},x_{j0}=x_{j}^{+},x_{j1}=x_{j}^{-}$. Let us see that $\widetilde{\gamma}$ is disjointness preserving. To do so, fix $i\in\{1,\dots,k\}$ and take $y_{i},\,z_{i}\in E_{i}$ such that $|y_{i}|\wedge|z_{i}|=0$. Given $x_{j}\in E_{j}^{+}$ for each $j\in\{1,\dots,k\}\setminus\{i\}$, we have
$|\widetilde{\gamma}(x_{1},\ldots,y_{i},\ldots,$ $x_{k})|\wedge|\widetilde{\gamma}(x_{1},\ldots,z_{i},\ldots,x_{k})|\leq$
$\leq\widetilde{\gamma}(x_{1},\ldots,|y_{i}|,\ldots,x_{k})\wedge\widetilde{\gamma}(x_{1},\ldots,|z_{i}|,\ldots,x_{k})$
$=\gamma(x_{1},\ldots,|y_{i}|,\ldots,x_{k})\wedge\gamma(x_{1},\ldots,|z_{i}|,\ldots,x_{k})$
$=A(x_{1},\ldots,|y_{i}|,\ldots,x_{k},x_{0}^{k+1})\wedge A(x_{1},\ldots,|z_{i}|,\ldots,x_{k},x_{0}^{k+1})=0,$
because $A$ is disjointness preserving. This proves that $\widetilde{\gamma}\colon E_{1}\times\cdots\times E_{k}\to\mathbb{R}$ is a disjointness preserving positive $k$-linear form. By the induction hypothesis, there exist positive disjointness preserving linear functionals $\varphi_{1}\in E_{1}^{\sim},\ldots,\varphi_{k}\in E_{k}^{\sim}$ such that $\widetilde{\gamma}(x_{1},\dots,x_{k})=\varphi_{1}(x_{1})\cdots\varphi_{k}(x_{k})$ for all $x_{1}\in E_{1},\dots,x_{k}\in E_{k}$. Finally, defining $\varphi_{k+1}=A_{x_{0}^{1},\dots,x_{0}^{k}}$, we have, for all $x_{1}\in E_{1},\ldots,x_{k+1}\in E_{k+1}$,
$A(x_{1},\ldots,x_{k},x_{k+1})$ $=\sum_{{n_{i}=0,1}\atop{1\leq i\leq k}}(-1)^{\sum n_{i}}A(x_{1n_{1}},\ldots,x_{kn_{k}},x_{k+1})$
$=\sum_{{n_{i}=0,1}\atop{1\leq i\leq k}}(-1)^{\sum n_{i}}A_{x_{1n_{1}},\ldots,x_{kn_{k}}}(x_{k+1})$
$=\sum_{{n_{i}=0,1}\atop{1\leq i\leq k}}(-1)^{\sum n_{i}}\gamma(x_{1n_{1}},\ldots,x_{kn_{k}})A_{x_{0}^{1},\ldots,x_{0}^{k}}(x_{k+1})$
$=\widetilde{\gamma}(x_{1},\ldots,x_{k})\varphi_{k+1}(x_{k+1})=\varphi_{1}(x_{1})\cdots\varphi_{k}(x_{k})\varphi_{k+1}(x_{k+1}),$
and we are done. ∎
@@ 3 Arens extensions of disjointness preserving multilinear operators
We shall use the notation from [7] to introduce the Arens extensions of multilinear operators. Let $m\in\mathbb{N}$ and $\rho\in S_{m}$ be given. For every $i=1,\ldots,m-1$, and $x_{\rho(i)}^{\prime\prime}\in E_{\rho(i)}^{\sim\sim}$, we consider the operator
$\overline{x_{\rho(i)}^{\prime\prime}}\colon\mathcal{L}_{b}(E_{\rho(i)},\ldots,E_{\rho(m)})\longrightarrow\mathcal{L}_{b}(E_{\rho(i+1)},\ldots,E_{\rho(m)})~{},~{}\overline{x_{\rho(i)}^{\prime\prime}}(B)=x_{\rho(i)}^{\prime\prime}\circ B^{i},$
@@
$B^{i}(x_{i+1},\ldots,x_{m})(x_{i})=B(x_{i},x_{i+1},\ldots,x_{m}),\text{ for all }x_{j}\in E_{\rho(j)},j=i,\ldots,m.$
In the case $i=m$, we define $\overline{x_{\rho(m)}^{\prime\prime}}\colon E_{\rho(m)}^{\sim}\longrightarrow\mathbb{R}$ by $\overline{x_{\rho(m)}^{\prime\prime}}=x_{\rho(m)}^{\prime\prime}$. The following lemma will be needed later.
Lemma 3.1.
For every $\rho\in S_{m}$, the operators $\overline{x_{\rho(1)}^{\prime\prime}},\dots,\overline{x_{\rho(m)}^{\prime\prime}}$ defined above are linear and order bounded. Moreover, if $0\leq x_{\rho(i)}^{\prime\prime}\in E_{i}^{\sim\sim}$ for some $i\in\{1,\dots,m\}$, then $\overline{x_{\rho(i)}^{\prime\prime}}$ is positive.
Proof.
It is clear that $\overline{x_{\rho(1)}^{\prime\prime}},\dots,\overline{x_{\rho(m)}^{\prime\prime}}$ are linear and that $\overline{x_{\rho(m)}^{\prime\prime}}$ is order bounded. Combining [17, Lemma 2.11] and [5, p. 15], we get $\mathcal{L}_{b}(E_{\rho(k)},\ldots,E_{\rho(m)})=\mathcal{L}_{r}(E_{\rho(k)},\ldots,E_{\rho(m)})$ for every $k=1,\ldots,m-1$. It follows from the proof of [6, Proposition 2.1] that $\overline{x_{\rho(i)}^{\prime\prime}}(B)$ is an order bounded operator for every $B\in\mathcal{L}_{b}(E_{\rho(i)},\ldots,E_{\rho(m)})$ and every $i=1,\dots,m-1$. The last statement follows immediately from [6, Proposition 2.1]. ∎
@@
The Arens extension of an order bounded $m$-linear operator $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ with respect to the permutation $\rho$ is defined as the $m$-linear operator $AR_{m}^{\rho}(A)\colon E_{1}^{\sim\sim}\times\cdots\times E_{m}^{\sim\sim}\longrightarrow F^{\sim\sim}$ given by
@@
Let us see that the extensions $AR_{m}^{\rho}(A)$ are always order bounded.
Lemma 3.2.
Let $E_{1},\ldots,E_{m},F$ be Riesz spaces. If $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ is an $m$-linear order bounded operator, then $AR_{m}^{\rho}(A)$ is order bounded for every $\rho\in S_{m}$.
Proof.
Given $x_{1}^{\prime\prime}\in(E_{1}^{\sim\sim})^{+},\dots,x_{m-1}^{\prime\prime}\in(E_{m-1}^{\sim\sim})^{+}$, we need to prove that the map $B\colon E_{m}^{\sim\sim}\to F^{\sim\sim}$, defined by $B(x_{m}^{\prime\prime})=AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\ldots,x_{m}^{\prime\prime})$, is order bounded. Indeed, if $O\subset E_{m}^{\sim\sim}$ is an order bounded set, then there exists $x^{\prime\prime},y^{\prime\prime}\in E_{m}^{\sim\sim}$ such that $x^{\prime\prime}\leq x_{m}^{\prime\prime}\leq y^{\prime\prime}$ for every $x_{m}^{\prime\prime}\in O$. Thus $|x_{m}^{\prime\prime}|\leq y^{\prime\prime}+2|x^{\prime\prime}|$ for every $x_{m}^{\prime\prime}\in O$ and
$|B(x_{m}^{\prime\prime})|$ $=|AR_{m}^{\rho}(A)(x_{1}^{\prime\prime},\ldots,x_{m-1}^{\prime\prime},x_{m}^{\prime\prime})|\leq|AR_{m}^{\rho}(A)|(x_{1}^{\prime\prime},\ldots,x_{m-1}^{\prime\prime},|x_{m}^{\prime\prime}|)$
$\leq|AR_{m}^{\rho}(A)|(x_{1}^{\prime\prime},\ldots,x_{m-1}^{\prime\prime},y^{\prime\prime}+2|x^{\prime\prime}|)$
$\leq|AR_{m}^{\rho}(A)|(x_{1}^{\prime\prime},\ldots,x_{m-1}^{\prime\prime},y^{\prime\prime})+2|AR_{m}^{\rho}(A)|(x_{1}^{\prime\prime},\ldots,x_{m-1}^{\prime\prime},|x^{\prime\prime}|),$
which implies that $B$ is an order bounded operator, proving that $AR_{m}^{\rho}(A)$ is order bounded in the last variable. The same argument can be used to prove its order boundedness in the other variables. ∎
Our following result is an immediate consequence of Proposition 2.5.
Lemma 3.3.
Let $E_{1},\ldots,E_{m},F$ be Riesz spaces with $F$ Archimedean and let $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ be an $m$-linear order bounded operator. Then, $A$ is disjointness preserving if and only if the linear operator $A_{j}\colon E_{j}\longrightarrow F,\,A_{j}(x_{j})=A(x_{1}\ldots,x_{m})$, is disjointness preserving for all $j\in\{1,\ldots,m\}$ and $x_{i}\in E_{i},\,i\neq j$.
@@
Let $E_{1},E_{2},F$ be Riesz spaces with $F$ Archimedean and let $A\colon E_{1}\times E_{2}\longrightarrow F$ be an order bounded bilinear operator. If $A$ is disjointness preserving, then for each $\rho\in S_{2}$ and $x_{1}\in E_{1},x_{2}\in E_{2}$, the operators
@@
hence $A^{\ast\ast\ast}(J_{E_{1}}(x_{1}),x_{2}^{\prime\prime})=A(x_{1},\cdot)^{\prime\prime}(x_{2}^{\prime\prime})$ for all $x_{1}\in E_{1}$ and $x_{2}^{\prime\prime}\in E_{2}^{\sim\sim}$. Since $A(x_{1},\cdot)$ is disjointness preserving by Lemma 3.3, so is its second adjoint $A(x_{1},\cdot)^{\prime\prime}$ by Theorem 2.1(a). Therefore, $A^{\ast\ast\ast}(J_{E_{1}}(x_{1}),\cdot)$ is disjointness preserving. The same argument shows that $AR_{2}^{id}(A)(J_{E_{1}}(x_{1}),\cdot)$ is disjointness preserving.
@@
This proves that $A^{\ast\ast\ast}(x_{1}^{\prime\prime},J_{E_{2}}(x_{2}))=A(\cdot,x_{2})^{\prime\prime}(x_{1}^{\prime\prime})$ for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim}$ and $x_{2}\in E_{2}$. Since $A(\cdot,x_{2})$ is disjointness preserving by Lemma 3.3, we get from Theorem 2.1(a) that its second adjoint is also disjointness preserving, which implies that $A^{\ast\ast\ast}(\cdot,J_{E_{2}}(x_{2}))$ is disjointness preserving as well. The same argument shows that $AR_{2}^{id}(A)(J_{E_{1}}(x_{1}),\cdot)$ is disjointness preserving. ∎
As in [6, p. 11], a map taking values in a Riesz space is said to have finite rank if the sublattice generated by its range is finite dimensional. Now, we prove our first main result, which shows that all Arens extensions of a finite rank order bounded disjointness preserving multilinear operator into an Archimedean Riesz space are also disjointness preserving.
@@
Let $E_{1},\ldots,E_{m},F$ be Riesz space with $F$ Archimedean and let $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ be a finite rank order bounded $m$-linear operator. If $A$ is disjointness preserving, then all Arens extensions $AR_{m}^{\rho}(A)$ of $A$, $\rho\in S_{m}$, are disjointness preserving.
@@
We split the proof into three steps. First we prove the case $F=\mathbb{R}$. Given an order bounded disjointness preserving $m$-linear form $A\colon E_{1}\times\cdots\times E_{m}\to\mathbb{R}$, we obtain from Proposition 2.5 and Theorem 2.6 that there are disjointness preserving linear functionals $\varphi_{1}\in E_{1}^{\sim},\ldots,\varphi_{m}\in E_{m}^{\sim}$ such that $|A|(x_{1},\ldots,x_{m})=\varphi_{1}(x_{1})\cdots\varphi_{m}(x_{m})$ for all $x_{1}\in E_{1},\ldots,x_{m}\in E_{m}$. By the proof of [6, Theorem 3.2],
@@
holds for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim},\ldots,x_{m}^{\prime\prime}\in E_{m}^{\sim\sim}$ and $\rho\in S_{m}$. Since $\varphi_{1},\dots,\varphi_{m}$ are disjointness preserving, their second adjoints $\varphi_{1}^{\prime\prime},\dots,\varphi_{1}^{\prime\prime}$ are disjointness preserving by Theorem 2.1(a). Then, $AR_{m}^{\rho}(|A|)$ is disjointness preserving. Using Lemma 3.1 and the fact that the composition of positive operators is a positive operator, it is easy to check that the correspondence
$A\in\mathcal{L}_{b}(E_{1},\ldots,E_{m};F)\mapsto AR_{m}^{\rho}(A)\in\mathcal{L}_{b}(E_{1}^{\sim\sim},\ldots,E_{m}^{\sim\sim};F^{\sim\sim})$
is a positive linear operator. So, $A\leq|A|$ implies $AR_{m}^{\rho}(A)\leq AR_{m}^{\rho}(|A|)$, from which it follows that $AR_{m}^{\rho}(A)$ is also disjointness preserving.
@@
for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim},\dots,x_{m}^{\prime\prime}\in E_{m}^{\sim\sim}$. Considering $\mathbb{R}^{n}$ with the coordinatewise order, implies that $AR_{m}^{\rho}(A)$ is disjointness preserving for every $\rho\in S_{m}$.
Finally, let $F$ be an Archimedean Riesz space $F$ and let $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ be a finite rank order bounded disjointness preserving operator. Calling $G$ be the finite-dimensional sublattice of $F$ generated by the range of $A$, by $i\colon G\longrightarrow F$ we denote the inclusion operator. It is clear that $A_{1}\colon E_{1}\times\cdots\times E_{m}\longrightarrow G,A_{1}(x_{1},\ldots,x_{m})=A(x_{1},\ldots,x_{m})$, is a disjointness preserving $m$-linear operator. Since $G$ is Archimedean, by [20, Corollary 1, p. 70] there exists a Riesz isomorphism $I\colon G\to\mathbb{R}^{n}$, where $n=\dim G$. Thus, $I\circ A_{1}$ is an $\mathbb{R}^{n}$-valued disjointness preserving multilineara operator, so, by the second part of this proof, we obtain that $AR_{m}^{\rho}(I\circ A_{1})$ is disjointness preserving. From [6, Remark 2.3] we have
@@
Now we proceed to the case of arbitrary order bounded multilinear operators. The following two results shall be helpful.
@@
Let $E_{1},\dots,E_{m},\,F$ be Riesz spaces, let $A\colon E_{1}\times\cdots\times E_{m}\to F$ be an order bounded disjointness preserving $m$-linear map, and let $\rho\in S_{m}$ be given. If $y^{\prime}\in F^{\sim}$ is disjointness preserving, then $J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A)$ is disjointness preserving and
@@
If $y^{\prime}\in F^{\sim}$ is disjointness preserving, then $y^{\prime}\circ A\colon E_{1}\times\cdots\times E_{m}\to\mathbb{R}$ is an order bounded $m$-linear disjointness preserving operator. By Theorem 3.5 we get that $AR_{m}^{\rho}(y^{\prime}\circ A)$ is also disjointness preserving. The proof of [6, Proposition 3.4] yields that $AR_{m}^{\rho}(y^{\prime}\circ A)=J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A)$ holds, proving that $J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A)$ is disjointness preserving. Calling on Proposition 2.5,
@@
for all $y^{\prime\prime}\in F^{\sim\sim}$. Combining (2) and (3), we have
@@
Let $E_{1},\ldots,E_{m},F$ be Banach lattices, let $A\in{\cal L}_{b}(E_{1},\ldots,E_{m};F)$ be a disjointness preserving operator and let $\rho\in S_{m}$ be given. Then:
@@
(b) Fix $i\in\{1,\ldots,m\}$ and let $w_{i}^{\ast\ast},z_{i}^{\ast\ast}\in E_{i}^{\ast\ast}$ be such that $|w_{i}^{\ast\ast}|\wedge|z_{i}^{\ast\ast}|=0$. Then
$(|AB_{m}^{\rho}(A)(x_{1}^{**},\ldots,w_{i}^{\ast\ast},\ldots,x_{m}^{**})|\wedge|AB_{m}^{\rho}(A)(x_{1}^{**},\ldots,z_{i}^{\ast\ast},\ldots,x_{m}^{**})|)(y^{*})=0$

### v2

@@ 3 Arens extensions of disjointness preserving multilinear operators
We shall use the notation from [7] to introduce the Arens extensions of multilinear operators. Let $m\in\mathbb{N}$ and $\rho\in S_{m}$ be given. For $i=1,\ldots,m-1$, and $x_{\rho(i)}^{\prime\prime}\in E_{\rho(i)}^{\sim\sim}$, consider the operator
$\overline{x_{\rho(i)}^{\prime\prime}}\colon\mathcal{L}_{r}(E_{\rho(i)},\ldots,E_{\rho(m)})\longrightarrow\mathcal{L}_{r}(E_{\rho(i+1)},\ldots,E_{\rho(m)})~{},~{}\overline{x_{\rho(i)}^{\prime\prime}}(B)=x_{\rho(i)}^{\prime\prime}\circ B^{i},$
@@
$B^{i}(x_{i+1},\ldots,x_{m})(x_{i})=B(x_{i},x_{i+1},\ldots,x_{m})\text{ for all}x_{j}\in E_{\rho(j)},j=i,\ldots,m.$
In the case $i=m$, we define $\overline{x_{\rho(m)}^{\prime\prime}}\colon E_{\rho(m)}^{\sim}\longrightarrow\mathbb{R}$ by $\overline{x_{\rho(m)}^{\prime\prime}}=x_{\rho(m)}^{\prime\prime}$.
@@
The Arens extension of a regular $m$-linear operator $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ with respect to the permutation $\rho$ is defined as the $m$-linear operator $AR_{m}^{\rho}(A)\colon E_{1}^{\sim\sim}\times\cdots\times E_{m}^{\sim\sim}\longrightarrow F^{\sim\sim}$ given by
@@
Let $E_{1},E_{2},F$ be Riesz spaces with $F$ Archimedean, and let $A\colon E_{1}\times E_{2}\longrightarrow F$ be a regular bilinear operator. If $A$ is disjointness preserving, then, for $\rho\in S_{2}$ and $x_{1}\in E_{1},x_{2}\in E_{2}$, the operators
@@
hence $A^{\ast\ast\ast}(J_{E_{1}}(x_{1}),x_{2}^{\prime\prime})=A(x_{1},\cdot)^{\prime\prime}(x_{2}^{\prime\prime})$ for all $x_{1}\in E_{1}$ and $x_{2}^{\prime\prime}\in E_{2}^{\sim\sim}$. Since $A(x_{1},\cdot)$ is disjointness preserving by assumption, so is its second adjoint $A(x_{1},\cdot)^{\prime\prime}$ by Theorem 2.1(1). Therefore, $A^{\ast\ast\ast}(J_{E_{1}}(x_{1}),\cdot)$ is disjointness preserving. The same argument shows that $AR_{2}^{id}(A)(J_{E_{1}}(x_{1}),\cdot)$ is disjointness preserving.
@@
This proves that $A^{\ast\ast\ast}(x_{1}^{\prime\prime},J_{E_{2}}(x_{2}))=A(\cdot,x_{2})^{\prime\prime}(x_{1}^{\prime\prime})$ for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim}$ and $x_{2}\in E_{2}$. Since $A(\cdot,x_{2})$ is disjointness preserving by assumption, we get from Theorem 2.1(1) that its second adjoint is also disjointness preserving, which implies that $A^{\ast\ast\ast}(\cdot,J_{E_{2}}(x_{2}))$ is disjointness preserving as well. The same argument shows that $AR_{2}^{id}(A)(J_{E_{1}}(x_{1}),\cdot)$ is disjointness preserving. ∎
As in [6, p. 11], a map taking values in a Riesz space is said to have finite rank if the sublattice generated by its range is finite dimensional. Now, we prove our first main result, which shows that all Arens extensions of a finite rank regular disjointness preserving multilinear operator into a Dedekind complete Riesz space are also disjointness preserving.
@@
Let $E_{1},\ldots,E_{m},F$ be Riesz space with $F$ Dedekind complete and let $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ be a finite rank regular $m$-linear operator. If $A$ is disjointness preserving, then all Arens extensions $AR_{m}^{\rho}(A)$ of $A$, $\rho\in S_{m}$, are disjointness preserving.
@@
We split the proof into three steps. First we prove the case $F=\mathbb{R}$. Given a regular disjointness preserving $m$-linear form $A\colon E_{1}\times\cdots\times E_{m}\to\mathbb{R}$, from Proposition 2.4 we know that $|A|$ is a scalar-valued Riesz multimorphism. By [16, Theorem 6] there are scalar-valued Riesz homomorphisms $\varphi_{1}\in E_{1}^{\sim},\ldots,\varphi_{m}\in E_{m}^{\sim}$ such that $|A|(x_{1},\ldots,x_{m})=\varphi_{1}(x_{1})\cdots\varphi_{m}(x_{m})$ for all $x_{1}\in E_{1},\ldots,x_{m}\in E_{m}$. By the proof of [6, Theorem 3.2],
@@
holds for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim},\ldots,x_{m}^{\prime\prime}\in E_{m}^{\sim\sim}$ and $\rho\in S_{m}$. Since $\varphi_{1},\dots,\varphi_{m}$ are disjointness preserving, their second adjoints $\varphi_{1}^{\prime\prime},\dots,\varphi_{1}^{\prime\prime}$ are disjointness preserving by Theorem 2.1(1). Then, $AR_{m}^{\rho}(|A|)$ is disjointness preserving. By [6, Theorem 2.2(c)] we know that the correspondence
$A\in\mathcal{L}_{r}(E_{1},\ldots,E_{m};F)\mapsto AR_{m}^{\rho}(A)\in\mathcal{L}_{r}(E_{1}^{\sim\sim},\ldots,E_{m}^{\sim\sim};F^{\sim\sim})$
is a positive linear operator. So, $A\leq|A|$ implies $AR_{m}^{\rho}(A)\leq AR_{m}^{\rho}(|A|)$, which, together with the disjointness preservation of $AR_{m}^{\rho}(|A|)$, gives that $AR_{m}^{\rho}(A)$ is also disjointness preserving.
@@
for all $x_{1}^{\prime\prime}\in E_{1}^{\sim\sim},\dots,x_{m}^{\prime\prime}\in E_{m}^{\sim\sim}$. As $\mathbb{R}^{n}$ is endowed with the coordinatewise order, it follows that $AR_{m}^{\rho}(A)$ is disjointness preserving for every $\rho\in S_{m}$.
Finally, let $F$ be a Dedekind complete Riesz space and let $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ be a finite rank disjointness preserving regular operator. Calling $G$ be the finite-dimensional sublattice of $F$ generated by the range of $A$, by $i\colon G\longrightarrow F$ we denote the inclusion operator. It is clear that $A_{1}\colon E_{1}\times\cdots\times E_{m}\longrightarrow G,A_{1}(x_{1},\ldots,x_{m})=A(x_{1},\ldots,x_{m})$, is a disjointness preserving $m$-linear operator. Since $G$ is Archimedean, by [24, Corollary 1, p. 70] there exists a Riesz isomorphism $I\colon G\to\mathbb{R}^{n}$, where $n=\dim G$ and $\mathbb{R}^{n}$ is endowed with the coordinatewise order. Thus, $I\circ A_{1}$ is an $\mathbb{R}^{n}$-valued disjointness preserving multilinear operator, so $AR_{m}^{\rho}(I\circ A_{1})$ is disjointness preserving by the second part of this proof. From [6, Remark 2.3] we have
@@
Now we proceed to the case of arbitrary regular multilinear operators. The following two results shall be helpful.
@@
Let $E_{1},\dots,E_{m},\,F$ be Riesz spaces, let $A\colon E_{1}\times\cdots\times E_{m}\to F$ be a regular disjointness preserving $m$-linear operator, and let $\rho\in S_{m}$ be given. If $y^{\prime}\in F^{\sim}$ is disjointness preserving, then $J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A)$ is disjointness preserving and
@@
Since $A$ is regular and disjointness preserving and $y^{\prime}\in F^{\sim}$ is disjointness preserving, $y^{\prime}\circ A\colon E_{1}\times\cdots\times E_{m}\to\mathbb{R}$ is a regular disjointness preserving $m$-linear operator. By the proof of Theorem 3.2 we get that $AR_{m}^{\rho}(y^{\prime}\circ A)$ is also disjointness preserving. The proof of [6, Proposition 3.4] yields that $AR_{m}^{\rho}(y^{\prime}\circ A)=J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A)$ holds, proving that $J_{F^{\sim}}(y^{\prime})\circ AR_{m}^{\rho}(A)$ is disjointness preserving and that
@@
for every $y^{\prime\prime}\in F^{\sim\sim}$. Combining (2) and (3), we have
@@
Let $E_{1},\ldots,E_{m},F$ be Banach lattices, let $A\in{\cal L}_{r}(E_{1},\ldots,E_{m};F)$ be a disjointness preserving operator, and let $\rho\in S_{m}$ be given. Then:
@@
(b) Fix $i\in\{1,\ldots,m\}$ and let $w_{i}^{\ast\ast},z_{i}^{\ast\ast}\in E_{i}^{\ast\ast}$ be such that $w_{i}^{\ast\ast}\perp z_{i}^{\ast\ast}$. Then
$AB_{m}^{\rho}(A)(x_{1}^{**},\ldots,w_{i}^{\ast\ast},\ldots,x_{m}^{**})(y^{*})\perp AB_{m}^{\rho}(A)(x_{1}^{**},\ldots,z_{i}^{\ast\ast},\ldots,x_{m}^{**})|)(y^{*})\mbox{ in }\mathbb{R}$